# Enhanced Expressivity in Graph Neural Networks with Lanczos-Based Linear Constraints

### v1

@@ Introduction
Efforts aimed at augmenting the expressiveness of MPNNs have pursued four main directions: aligning with the $k$-WL hierarchy [34, 31, 4], enriching node features with identifiers, exploiting structural information that cannot be captured by the WL test [9, 8], and Subgraph GNNs (SGNNs) [6, 20]. SGNNs are a recent class of expressive GNNs that model graphs through a collection of subgraphs. They have emerged as a potential solution, extending GNNs by collecting extracted subgraphs explicitly or implicitly and incorporating these into the GNN architecture. Subgraph extraction can be achieved, e.g., by removing or marking specific nodes or directly counting specific substructures [38]. However, in the worst case, previous SGNNs involve computationally intensive preprocessing steps or running a GNN many times.
@@
Figure 1: The comprehensive workflow of LLwLC: We introduce a novel eigenbasis to boost the expressivity of GNNs by satisfying constraints $\mathbf{C}$. For the two $2$-regular input graphs (first dotted block), the constraint matrix $\mathbf{C}$ (two columns shown in the second dotted block) is derived. Specifically, as depicted, for the first Neumann eigenvalue constraint (focused on link representation), we integrate the degrees of nodes into the constraint matrix such that the summation over adjacent nodes in the boundary $x\sim y$ ensures $\bm{f}(x)-\bm{f}(y)=0$. To predict the dotted link, nodes up to two hops away are analyzed: one-hop in light blue and two-hop in dark blue, centered on yellow query nodes. Based on the input graph Laplacian matrix $\mathbf{L}$ and the constraint matrix $\mathbf{C}$, we address the eigenvalue problem under these linear constraints using the Lanczos algorithm with the linear constraint consisting of an outer loop (Lanczos algorithm) and an inner loop ($\mathbf{P}(\nu)$ to $\mathcal{N}(\mathbf{C})$ with LSQR algorithm) resulting in the eigenpairs $\mathbf{V}=\mathbf{Q}\mathbf{B}$ and $\mathbf{R}$. The third block demonstrates constructing eigenbases for the two 2-regular graphs, ensuring features conform to constraints (indicated in blue). We stack multiple blocks, where each learns the new features by applying a multi-layer perceptron ($f$) over the eigenvalue matrix $\mathbf{R}$, so we have $\sigma(\mathbf{V}f(\mathbf{R})\mathbf{V}{{}^{\top}}\mathbf{X}\mathbf{W})$. This sequence concludes with global pooling and a fully connected layer to output the predicted link probability $\hat{y}$.
The LLwLC eigenbasis enhances feature expressiveness by incorporating linear constraints derived from the graph structure. We propose two novel subgraph extraction policies, focusing on vertex-deleted subgraphs and Neumann eigenvalue constraints. We provide a conjecture that vertex-deleted subgraphs have the universal approximator capabilities of LLwLC, while Neumann constraints enable the encoding of boundary conditions and link representations between nodes, allowing to differentiate $k$-regular graphs that are indistinguishable by WL. Theoretical analysis indicates that LLwLC can be applied in various problem settings. We evaluate its effectiveness in link prediction tasks, where expressivity and particularly the ability to distinguish automorphic nodes is pivotal.
@@
A graph $G(V,E,\mathbf{X})$ consists of a vertex set $V$, edge set $E$, and node features $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $n$ is the number of nodes. Each column $\bm{x}_{v}\in\mathbb{R}^{d}$ represents the features of node $v\in V$. $\mathbf{A}$ and $\mathbf{D}$ are the graph’s adjacency and degree matrices, respectively. The graph Laplacian is $\mathbf{L}=\mathbf{D}-\mathbf{A}$, with $\mathbf{U}$ and $\mathbf{\Lambda}$ as its eigenvector and eigenvalue matrices. The degree of node $v$ is denoted by $d_{v}$.
@@
For a graph signal $\bm{x}\in\mathbb{R}^{n}$[43] defines the graph Fourier transform, $\mathbf{U}^{\top}\bm{x}$, and its inverse, $\mathbf{U}\bm{x}$, based on the eigenbasis of the graph Laplacian matrix $\mathbf{L}$. The graph convolution is $\mathbf{U}(\mathbf{U}^{\top}\bm{x}\ast\mathbf{U}^{\top}\bm{y})=\mathbf{U}g(\mathbf{\Lambda})\mathbf{U}^{\top}\bm{x}$ where $\bm{y}$ is the graph filter, $g$ is the function applied over the eigenvalue matrix $\mathbf{\Lambda}$ to encode the graph filter, and $\ast$ is the elementwise multiplication. The seminal spectral GCN method [12] is cubic in the number of nodes. To address this, different $g$ functions were defined [23, 3]. LanczosNet [29] uses the Lanczos algorithm for fast multi-scale computation with learnable spectral filters. However, like previous GNNs, LanczosNet has limited expressivity, e.g., it cannot distinguish between the $k$-regular graphs. To address these limitations and enhance feature expressiveness in GNNs, we introduce a novel learnable spectral basis for encoding subgraphs as linear constraints.
Lanczos with Linear Constraint Networks
@@
1:Input: $\mathbf{L},\mathbf{P}=\mathbf{I}-\mathbf{C}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}$, random initialized vector $\bm{\nu},$ $\kappa-$step LLwLC algorithm, error tolerance $\epsilon$
@@
16:return $\mathbf{V}=\mathbf{Q}\cdot\mathbf{B}$ and $\mathbf{R}$
@@
2:Init: $Y_{0}=X$
3:for $j=0$ to $k-1$ do
4:     $\mathbf{Y}_{i+1}=\sigma(\mathbf{V}f_{i}(\mathbf{R})\mathbf{V}^{\top}\mathbf{Y}_{i}\mathbf{W}_{i})$
5:end for
6:$\hat{y}$ = Fully Connected(Global Pooling($\mathbf{Y}_{k-1}$))
7:return $\hat{y}$
Lanczos Algorithm with Linear Constraints
For a given symmetric matrix $\mathbf{L}\in\mathbb{R}^{n\times n}$ and a randomly initialized vector $\bm{\nu}\in\mathbb{R}^{n}$, the $\kappa$-step Lanczos algorithm [27] computes an orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{n\times m}$ and a symmetric tridiagonal matrix $\mathbf{T}\in\mathbb{R}^{m\times m}$, such that $\mathbf{Q}^{\top}\mathbf{L}\mathbf{Q}=\mathbf{T}$. We represent $\mathbf{Q}_{N}=[\bm{q}_{1},\dots,\bm{q}_{N}]$ where the column vector $\bm{q}_{i}$ corresponds to the $i^{th}$ Lanczos vector. The matrices $\mathbf{B}\in\mathbb{R}^{m\times m}$ and $\mathbf{R}\in\mathbb{R}^{m\times m}$ represent the eigenvectors and eigenvalues of $\mathbf{T}$, respectively. By investigating the $j^{th}$ column of the system $\mathbf{L}\mathbf{Q}=\mathbf{Q}\mathbf{T}$ and rearranging terms, we obtain $\mathbf{L}\bm{q}_{j}=\beta_{j+1}\bm{q}_{j+1}+\beta_{j}\bm{q}_{j-1}+\alpha_{j}\bm{q}_{j}$.
@@
The construction of the constraint matrix, denoted as $\mathbf{C}$, is based on subgraphs extracted from the input graph. We address the creation of $\mathbf{C}$ through two distinct approaches: induced subgraphs and link representation. For the induced subgraph case, we populate its corresponding column in $\mathbf{C}$ with the degrees of each node (and fill the remaining entries of the column with 00 for nodes not involved in the subgraph), ensuring that the derived eigenvectors fulfill the condition $\sum\bm{f}(x)d_{x}=0$. This condition ensures that the features learned using this eigenbasis reflect the imposed constraints. In the case of link representation, we design the constraint matrix to ensure that the aggregate of eigenvector differences across nodes equates to zero, as expressed by $\sum(\bm{f}(x)-\bm{f}(y))=0$. The column corresponding to the subgraph is constructed based on the equation. The degrees of the nodes involved in the constraint are entered into the column to satisfy the equation. Specifically, the degrees of nodes two-hop-away are negated, and the degrees of nodes one-hop-away are included only if they are connected to nodes two hops away. The remaining entries for nodes not involved are filled with 00 (see Figure 1 for an example). We build the other columns of the constraint matrix accordingly. A single matrix encompasses all constraints and each column representing a distinct constraint (${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}{\sum}{}{}\bm{f}(x)d_{x}=0}$, $[{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}]$ ). Implementing these constraints and requiring the graph Laplacian matrix’s eigenvectors to comply with them allows for feature extraction aligned with the eigenbasis constraints. This enables nuanced differentiation between structures, such as $k$-regular graphs, as shown in Figure 1.
Addressing the eigenvalue problem with linear constraints yields a tridiagonal matrix, denoted as $\mathbf{T}$, and an orthogonal matrix, $\mathbf{Q}$. The decomposition of matrix $\mathbf{T}$ produces matrices $\mathbf{R}$ and $\mathbf{B}$. Here, $\mathbf{R}$ represents the Ritz eigenvalues, and $\mathbf{V}=\mathbf{Q}\mathbf{B}$ forms the eigenbasis that satisfies the constraints imposed by matrix $\mathbf{C}$. Forcing the eigenvectors to satisfy the constraints leads to having different eigenbasis for graphs where the MPNN returns the same features. This is exemplified in the case described in Figure 1, where two 2-regular graphs yield two different eigenbases.
@@
After determining the eigenbasis, we are ready to establish the full block of the Lanczos Layer with Linear Constraint (LLwLC). In this new eigenbasis, we develop spectral filters by applying a multilayer perceptron ($f$) to the eigenvalue matrix $\mathbf{R}$. With these learned filters, we reconstruct our basis and transform the graph signals $\mathbf{X}\in\mathbb{R}^{m\times n}$ into this basis to extract features that meet our specific constraints. $\hat{\mathbf{L}}$ is the graph Laplacian matrix computed from the low-rank approximation of the constrained eigenvalue problem. Each LLwLCNet block is
@@
Lanczos Algorithm with Linear Constraint Convergence
@@
The accuracy of the linear least square problem using QR factorization depends on the precision of the QR factorization. As discussed by  Zhang, Baharlouei, and Wu [56], two types of accuracy errors are crucial in QR factorization when solving linear least square problems: The backward error for a matrix $\mathbf{Z}$ is defined as $\frac{\lVert\mathbf{Z}-\hat{\mathbf{Q}}\hat{\mathbf{R}}\rVert}{\lVert\mathbf{Z}\rVert}$ and the orthogonality error of $\hat{\mathbf{Q}}$ is measured by $\lVert\mathbf{I}-\hat{\mathbf{Q}}{{}^{\top}}\hat{\mathbf{Q}}\rVert$. Ideally, both numerical errors should be zero, but due to roundoff errors and the potential loss of orthogonality in the Gram-Schmidt QR process, the QR factorization might not be sufficiently accurate for solving the linear least square problem.
After examining the impact on accuracy, we analyze the theoretical gap between the exact Lanczos algorithm and its perturbed variant due to inexact QR factorization. The inexact QR factorization applied to solve 2 will impact the accuracy of both the Lanczos vectors and the tridiagonal matrices produced. Consequently, the computed tridiagonal matrix $\mathbf{T}_{j}$ is a perturbed version of the theoretical tridiagonal matrix, denoted as $\mathbf{T}^{\ast}_{j}$, that would be generated by an exact Lanczos iteration. This relationship can be expressed as $\mathbf{T}_{j}=\mathbf{T}^{\ast}_{j}+\mathbf{E}_{j}$, where $\mathbf{E}_{j}$ is the perturbation matrix after the $j^{th}$ step. The following theorem details the error bounds of the perturbed tridiagonal matrix in comparison to the theoretical exact solution of $\mathbf{T}$ after the $j^{th}$ step of the Lanczos algorithm.
@@
The tridiagonal matrix $\mathbf{T}_{j}$ generated at the end of the $j^{th}$ finite precision Lanczos process satisfying $\mathbf{L}\mathbf{Q}_{j}=\mathbf{Q}_{j}\mathbf{T}_{j}+\beta_{j+1}\bm{q}_{j+1}\bm{e}{{}^{\top}}_{j}+\mathbf{F}_{j},$ where $\bm{e}_{j}^{\top}$ is a vector with the $j^{th}$ component one and all the other components zero, $\mathbf{F}=(\bm{f}_{1},\dots,\bm{f}_{j})$ is the perturbation term with $\lVert\bm{f}_{j}\rVert_{2}\leq\epsilon\lVert\mathbf{L}\rVert_{2},\epsilon\ll 1$, is the same as that generated by an exact Lanczos process but with a different matrix $\tilde{\mathbf{L}}$. The matrices $\mathbf{L}$ and $\tilde{\mathbf{L}}$ are close in the sense that for any eigenvalue $\lambda(\tilde{\mathbf{L}})$ of $\tilde{\mathbf{L}}$, there is an eigenvalue $\lambda(\mathbf{L})$ of $\mathbf{L}$ such that $|\lambda(\tilde{\mathbf{L}})-\lambda({\mathbf{L}})|\leq\lVert\mathbf{F}_{j}\rVert_{2}.$ Therefore, in our case with the constant accuracy of the QR factorization, we can show $\mathbf{P}\mathbf{L}\mathbf{P}\tilde{\mathbf{Q}}_{j}=\tilde{\mathbf{Q}}_{j}\mathbf{T}_{j}+\beta_{j}\tilde{\bm{q}}_{j+1}\bm{e}_{j}^{\top}+\tilde{\mathbf{F}}_{j},$ where $\tilde{\mathbf{F}}_{j}=\mathbf{O}(\eta)$ with $\eta$ corresponds to the accuracy of the QR method.
Having established each perturbed Lanczos corresponds to an exact algorithm, we demonstrate the theorem below to bound the approximation error, as discussed in [29].
@@
Based on Greenbaum’s results [19], for our computed perturbed Lanczos algorithm, exists an exact Lanczos algorithm but for a different matrix. Based on 2, we also cognize the upper bound of the low-rank approximator of the Lanczos algorithm. Thus, the perturbed Lanczos algorithm, caused by the inaccuracy of the QR method for solving the least square equation, converges to the upper bound of the low-rank approximation of the matrix of the exact Lanczos algorithm.
Subgraph Extraction Policy

### v2

# Boosting Graph Neural Network Expressivity with Learnable Lanczos Constraints

@@ 1 Introduction
Efforts aimed at augmenting the expressiveness of MPNNs have pursued four main directions: aligning with the $k$-WL hierarchy [33, 31, 4], enriching node features with identifiers, exploiting structural information that cannot be captured by the WL test [9, 8], and Subgraph GNNs (SGNNs) [6, 20]. SGNNs are a recent class of expressive GNNs that model graphs through a collection of subgraphs, extracted explicitly or implicitly. Subgraph extraction can be achieved, e.g., by removing or marking specific nodes or directly counting specific substructures [38]. However, in the worst case, previous SGNNs involve computationally intensive preprocessing steps or running a GNN many times.
@@
Figure 1: The comprehensive workflow of LLwLC: We introduce a novel eigenbasis to boost the expressivity of GNNs by introducing constraints. For the two $2$-regular input graphs (first dotted block), the constraint matrix $\mathbf{C}$ (two columns shown in the second dotted block) is derived. Specifically, as depicted, for the first Neumann eigenvalue constraint (focused on link representation), we integrate the degrees of nodes into one column of the constraint matrix $\bm{C}$ such that $\mathbf{C}{{}^{\top}}\bm{f}=0$ ensures $\bm{f}(x)-\bm{f}(y)=0$ for adjacent nodes $x$ and $y$ in the boundary. To predict the dotted link, nodes up to two hops away are analyzed: one-hop in light blue and two-hop in dark blue, centered on yellow query nodes. Based on the input graph Laplacian matrix $\mathbf{L}$ and the constraint matrix $\mathbf{C}$, we address the eigenvalue problem under these linear constraints using the Lanczos algorithm with the linear constraint consisting of an outer loop (Lanczos algorithm) and an inner loop ($\mathbf{P}(\nu)$ to $\mathcal{N}(\mathbf{C})$ with LSQR algorithm) resulting in the eigenpairs $\mathbf{V}=\mathbf{Q}\mathbf{B}$ and $\mathbf{R}$. The third block demonstrates constructing eigenbases for the two 2-regular graphs, ensuring features conform to constraints (indicated in blue). We stack multiple blocks, where block $i$ learns the new features $\mathbf{X}_{i+1}=\sigma(\mathbf{V}f_{i}(\mathbf{R})\mathbf{V}{{}^{\top}}\mathbf{X}_{i}\mathbf{W})$ by applying a multi-layer perceptron $f_{i}$ to the eigenvalue matrix $\mathbf{R}$. This sequence concludes with global pooling and a fully connected layer to output the predicted link probability $\hat{y}$.
The LLwLC eigenbasis enhances feature expressiveness by incorporating linear constraints derived from the graph structure. We propose two novel subgraph extraction policies, focusing on vertex-deleted subgraphs and Neumann eigenvalue constraints. By leveraging vertex-deleted subgraphs, LLwLC can effectively distinguish between graphs that are not separable using the 2-WL test, while Neumann constraints enable the encoding of boundary conditions and link representations between nodes, allowing differentiation of $k$-regular graphs that are indistinguishable by the WL test. Theoretical analysis indicates that LLwLC can be applied in various problem settings. We evaluate its effectiveness in link prediction tasks, where expressivity and particularly the ability to distinguish automorphic nodes is pivotal.
@@ 2 Preliminaries
An undirected graph $G(V,E,\mathbf{X})$ consists of a vertex set $V$, edge set $E$, and node features $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $n$ is the number of nodes. Each row $\bm{x}_{v}\in\mathbb{R}^{d}$ represents the features of node $v\in V$. $\mathbf{A}$ and $\mathbf{D}$ are the graph’s adjacency and degree matrices, respectively. The graph Laplacian is $\mathbf{L}=\mathbf{D}-\mathbf{A}$, with $\mathbf{U}$ and $\mathbf{\Lambda}$ as its eigenvector and eigenvalue matrices. The degree of node $v$ is denoted by $d_{v}$.
@@
For a graph signal $\bm{x}\in\mathbb{R}^{n}$human et al. [43] define the graph Fourier transform, $\mathbf{U}^{\top}\bm{x}$, and its inverse, $\mathbf{U}\bm{x}$, based on the eigenbasis of the graph Laplacian matrix $\mathbf{L}$. The graph convolution is $\mathbf{U}(\mathbf{U}^{\top}\bm{x}\ast\mathbf{U}^{\top}\bm{y})=\mathbf{U}g(\mathbf{\Lambda})\mathbf{U}^{\top}\bm{x}$ where $\bm{y}$ is the graph filter, $g$ is the function applied over the eigenvalue matrix $\mathbf{\Lambda}$ to encode the graph filter, and $\ast$ is the elementwise multiplication. The seminal spectral GCN method [12] is cubic in the number of nodes. To address this, different $g$ functions were defined [23, 3]. LanczosNet [29] uses the Lanczos algorithm for fast multi-scale computation with learnable spectral filters. However, like previous GNNs, LanczosNet has limited expressivity, e.g., it cannot distinguish between the $k$-regular graphs. To address these limitations and enhance feature expressiveness in GNNs, we introduce a novel learnable spectral basis for encoding subgraphs as linear constraints.
3 Lanczos with Linear Constraint Networks
@@
1:  Input: $\mathbf{L},\mathbf{P}=\mathbf{I}-\mathbf{C}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}$, random initialized vector $\bm{\nu},$ number of steps $\kappa$, error tolerance $\epsilon$
@@
16:  Return $\mathbf{V}=\mathbf{Q}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\cdot}\mathbf{B}$ and $\mathbf{R}$
@@
2:  Init: $\mathbf{X}_{0}=\mathbf{X}$
3:  for $i=0$ to $k-1$ do
4:    $\mathbf{X}_{i+1}=\sigma(\mathbf{V}f_{i}(\mathbf{R})\mathbf{V}^{\top}\mathbf{X}_{i}\mathbf{W}_{i})$
[…]
6:  Return $\hat{y}$
3.1 Lanczos Algorithm with Linear Constraints
For a given symmetric matrix $\mathbf{L}\in\mathbb{R}^{n\times n}$ and a randomly initialized vector $\bm{\nu}\in\mathbb{R}^{n}$, the $\kappa$-step Lanczos algorithm [27] computes an orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{n\times m}$ and a symmetric tridiagonal matrix $\mathbf{T}\in\mathbb{R}^{m\times m}$, such that $\mathbf{Q}^{\top}\mathbf{L}\mathbf{Q}=\mathbf{T}$. We represent $\mathbf{Q}_{N}=[\bm{q}_{1},\dots,\bm{q}_{N}]$ where the column vector $\bm{q}_{i}$ corresponds to the $i$th Lanczos vector. The matrices $\mathbf{B}\in\mathbb{R}^{m\times m}$ and $\mathbf{R}\in\mathbb{R}^{m\times m}$ represent the eigenvectors and eigenvalues of $\mathbf{T}$, respectively. By investigating the $j$th column of the system $\mathbf{L}\mathbf{Q}=\mathbf{Q}\mathbf{T}$ and rearranging terms, we obtain $\mathbf{L}\bm{q}_{j}=\beta_{j+1}\bm{q}_{j+1}+\beta_{j}\bm{q}_{j-1}+\alpha_{j}\bm{q}_{j}$.
@@ 3.2 LLwLCNet
The matrix $\mathbf{C}$ allows us to specify linear constraints that the graph Laplacian matrix’s eigenvectors must satisfy. Each column of the constraint matrix represents a distinct constraint. We introduce two approaches for defining the constraint matrix $\mathbf{C}$ on the basis of specific subgraphs: vertex-deleted subgraphs and Neumann eigenvalue constraints for link representations. A detailed explanation of how each column of the constraint matrix is built for these methods can be found in Section 3.4.
Addressing the eigenvalue problem with linear constraints yields a tridiagonal matrix, denoted as $\mathbf{T}$, and an orthogonal matrix $\mathbf{Q}$. The decomposition of matrix $\mathbf{T}$ produces matrices $\mathbf{R}$ and $\mathbf{B}$. Here, $\mathbf{R}$ represents the Ritz eigenvalues, and $\mathbf{V}=\mathbf{Q}\mathbf{B}$ forms the eigenbasis that satisfies the constraints imposed by matrix $\mathbf{C}$. Forcing the eigenvectors to satisfy appropriately defined constraints leads to having different eigenbasis for graphs where the MPNN returns the same features. This is exemplified in the case described in Figure 1, where two $2$-regular graphs yield two different eigenbases.
@@
After determining the eigenbasis, we are ready to establish the full block of the Lanczos Layer with Linear Constraint (LLwLC). In this new eigenbasis, we develop spectral filters by applying a multilayer perceptron $f$ to the eigenvalue matrix $\mathbf{R}$. With these learned filters, we reconstruct our basis and transform the graph signals $\mathbf{X}\in\mathbb{R}^{m\times n}$ into this basis to extract features that meet our specific constraints. $\hat{\mathbf{L}}$ is the graph Laplacian matrix computed from the low-rank approximation of the constrained eigenvalue problem. Each LLwLCNet block is
@@
3.3 Lanczos Algorithm with Linear Constraint Convergence
@@
The accuracy of the linear least square problem using QR factorization depends on the precision of the QR factorization. As discussed by  Zhang et al. [56], two types of accuracy errors are crucial in QR factorization when solving linear least square problems: The backward error for a matrix $\mathbf{Z}$ is defined as $\frac{\lVert\mathbf{Z}-\hat{\mathbf{Q}}\hat{\mathbf{R}}\rVert}{\lVert\mathbf{Z}\rVert}$ and the orthogonality error of $\hat{\mathbf{Q}}$ is measured by $\lVert\mathbf{I}-\hat{\mathbf{Q}}{{}^{\top}}\hat{\mathbf{Q}}\rVert$. Ideally, both numerical errors should be zero, but due to roundoff errors and the potential loss of orthogonality in the Gram-Schmidt QR process, the QR factorization might not be sufficiently accurate for solving the linear least square problem.
After examining the impact on accuracy, we analyze the theoretical gap between the exact Lanczos algorithm and its perturbed variant due to inexact QR factorization. The inexact QR factorization applied to solve Eq. (2) will impact the accuracy of both the Lanczos vectors and the tridiagonal matrices produced. Consequently, the computed tridiagonal matrix $\mathbf{T}_{j}$ is a perturbed version of the theoretical tridiagonal matrix, denoted as $\mathbf{T}^{\ast}_{j}$, that would be generated by an exact Lanczos iteration. This relationship can be expressed as $\mathbf{T}_{j}=\mathbf{T}^{\ast}_{j}+\mathbf{E}_{j}$, where $\mathbf{E}_{j}$ is the perturbation matrix after the $j^{th}$ step. The following theorem details the error bounds of the perturbed tridiagonal matrix in comparison to the theoretical exact solution of $\mathbf{T}$ after the $j^{th}$ step of the Lanczos algorithm.
@@
The tridiagonal matrix $\mathbf{T}_{j}$ generated at the end of the $j$th finite precision Lanczos process satisfying $\mathbf{L}\mathbf{Q}_{j}=\mathbf{Q}_{j}\mathbf{T}_{j}+\beta_{j+1}\bm{q}_{j+1}\bm{e}{{}^{\top}}_{j}+\mathbf{F}_{j},$ where $\bm{e}_{j}^{\top}$ is a vector with the $j$th component one and all the other components zero, $\mathbf{F}=(\bm{f}_{1},\dots,\bm{f}_{j})$ is the perturbation term with $\lVert\bm{f}_{j}\rVert_{2}\leq\epsilon\lVert\mathbf{L}\rVert_{2},\epsilon\ll 1$, is the same as that generated by an exact Lanczos process but with a different matrix $\tilde{\mathbf{L}}$. The matrices $\mathbf{L}$ and $\tilde{\mathbf{L}}$ are close in the sense that for any eigenvalue $\lambda(\tilde{\mathbf{L}})$ of $\tilde{\mathbf{L}}$, there is an eigenvalue $\lambda(\mathbf{L})$ of $\mathbf{L}$ such that $|\lambda(\tilde{\mathbf{L}})-\lambda({\mathbf{L}})|\leq\lVert\mathbf{F}_{j}\rVert_{2}.$ Therefore, in our case with the constant accuracy of the QR factorization, we can show $\mathbf{P}\mathbf{L}\mathbf{P}\tilde{\mathbf{Q}}_{j}=\tilde{\mathbf{Q}}_{j}\mathbf{T}_{j}+\beta_{j}\tilde{\bm{q}}_{j+1}\bm{e}_{j}^{\top}+\tilde{\mathbf{F}}_{j},$ where $\tilde{\mathbf{F}}_{j}=\mathbf{O}(\eta)$ with $\eta$ corresponds to the accuracy of the QR method.
Having established each perturbed Lanczos algorithm corresponds to an exact Lanczos algorithm, we demonstrate the theorem below to bound the approximation error, as discussed in [29].
@@
Based on Greenbaum’s results [19], for our computed perturbed Lanczos algorithm, exists an exact Lanczos algorithm but for a different matrix. Based on Theorem 2, we also cognize the upper bound of the low-rank approximator of the Lanczos algorithm. Thus, the perturbed Lanczos algorithm, caused by the inaccuracy of the QR method for solving the least square equation, converges to the upper bound of the low-rank approximation of the matrix of the exact Lanczos algorithm.
3.4 Subgraph Extraction Policy